# Corner modes in Non-Hermitian long-range model

### v1

@@ III.2 NH Model with second NN hopping
We now examine the energy spectrum for the second NN model in Figs. 3(d,e,f) employing OBC. Similar to the first NN case, the momentum takes the following non-Bloch form $k_{i}\to k^{\prime}_{i}-i\gamma_{2}$ with $\gamma_{2}=\gamma/(\lambda^{s}_{1}+2\lambda^{s}_{2})$ and $i=x,y$. This leads to the renormalized mass term $m^{\prime}_{0}=m_{0}-4-5\gamma_{2}^{2}<0~{}(>0)$ for the topological (trivial) phase with zero-energy (finite energy bulk) modes considering the Bloch momentum $\boldsymbol{\mathrm{k}}=(0,0)$. On the other hand, another topological (trivial) phase with zero-energy (finite energy) modes appears for $-m^{\prime}_{0}=m_{0}+2+\gamma_{2}^{2}>0~{}(<0)$ while exploiting energy around the Bloch momentum $\boldsymbol{\mathrm{k}}=(\pm 2\pi/3,\pm 2\pi/3)$. Our analysis indicates the existence of four [sixteen] corner modes for $3\gamma_{2}^{2}<m_{0}<4+5\gamma_{2}^{2}$ [$-2-\gamma^{2}_{2}<m_{0}<3\gamma_{2}^{2}$] yielding ${\rm Re}[E]=0$ under OBC. However, the numerical findings shown in Fig. 3(d) and the topological regime highlighted by the red lines, do not fully match with the exceptional boundaries predicted analytically. This can be attributed to the finite size effect in the second NN case, which is substantially small for the first NN case. We find macroscopic degeneracies at ${\rm Re}[E]=0$ around $m_{0}=0$, unlike the previous case. The complex energy spectrum ${\rm Im}[E]$ does not manifest any noteworthy features, as shown in Figs. 3(b,e), irrespective of PBC and OBC cases. The absolute value of energy is expected to vanish i.e.,  $|E|=0$ under OBC for $-2-\gamma^{2}_{2}<m_{0}<4+5\gamma_{2}^{2}$. However, numerical results suffer from finite size effects as far as the exceptional boundaries are concerned [see the red lines in Fig. 3(f)]. The finite value of $\gamma$ (NH effect) thus extends the Hermitian topological phase beyond its boundaries, leading to exceptional topological phases.
@@ III.3 Quadrupole Winding number
We now examine the present situation with $\gamma\neq 0$ where the NH analog of QWN is discussed. Importantly, CS is also preserved for the NH Hamiltonian $CH_{\gamma}(\boldsymbol{\mathrm{k}})C^{-1}=-H_{\gamma}(\boldsymbol{\mathrm{k}})$ allowing for the anti-diagonal form of $\tilde{H}_{\gamma}$. At the same time, the definition of $U^{A,B}_{C}$ remains unaltered for the NH case. We adopt the bi-orthogonalized definitions of $U_{A}$ and $U^{\dagger}_{B}$, obtained from the singular value decomposition of $h$, to define $U^{\dagger}_{A}$ and $U_{B}$, respectively. One has to ensure $\sum_{n}|U^{R}_{\sigma,n}\rangle\langle U^{L}_{\sigma,n}|=\mathds{1}$ and $\langle U^{L}_{\sigma,n}|U^{R}_{\sigma,m}\rangle=\delta_{mn}$ with $\sigma=A,B$ and $L(R)$ denotes the left (right) eigenvectors. This results in left [right] eigenvector corresponding to right eigenvector $(U^{\dagger}_{B})^{\dagger}\equiv U^{R}_{B}$ [left eigenvector $U^{\dagger}_{A}\equiv U^{L}_{A}$] as $U^{\dagger}_{B}\equiv U^{L}_{B}$ [$U_{A}\equiv U^{R}_{A}$]. Therefore, sublattice quadrupole operator $\mathcal{Q}_{\sigma}$ takes the form $\mathcal{Q}_{\sigma}=U_{\sigma}^{L}U^{\sigma}_{C}\mathcal{Q}U^{\sigma}_{C}U^{R}_{\sigma}$. In addition, the non-Bloch form of momentum has to be incorporated while computing $\mathcal{Q}_{\sigma}$.
@@
To be precise, the complex momentum $k_{i}\to k^{\prime}_{i}-i\gamma_{2}$ with $i=x,y$ leads to the exponentially enhanced and suppressed hopping elements by the multiplicative factors $\exp(\gamma_{1})$ and $\exp(-\gamma_{1})$ [$\exp(\gamma_{2})$ and $\exp(-\gamma_{2})$] for first [second] NN models. We use the real space form of the tight-binding model with the renormalized hopping amplitudes as follows $\lambda^{s,h,f}_{1,2}\to\lambda^{s,h,f}_{1,2}\exp(\gamma_{1,2})$ and $\lambda^{s,h,f}_{1,2}\to\lambda^{s,h,f}_{1,2}\exp(-\gamma_{1,2})$ for forward and backward hopping amplitudes, respectively. We consider the real space version of NH Hamiltonian $H_{\gamma}$ with the above mentioned renormalized hoppings in order to compute QWN. Altogether this enables us to define the NH analog of QWN $N_{xy}$ with dressed hopping and bi-orthogonalized definition. Note that, the real part of $N_{xy}$ exhibits quantized value for the present case with $\gamma\neq 0$ as demonstrated below.
@@
We now discuss the phase diagram of the first and second NN NH model in the $\gamma$-$m_{0}$ plane in Figs. 5(a,b) respectively. Note that, there exists four corner modes with ${\rm Re}[E]=0$ for $m_{0}<2+\gamma_{1}^{2}$, yielding $N_{xy}=1$. This refers to the fact that there exists only one topological zero-mode per corner [see Fig. 5(a)]. On the other hand, when $m_{0}>2+\gamma_{1}^{2}$, the NH model does not host any topological phase and hence $N_{xy}=0$. Hence, the topological phase boundary is $m_{0}=2+\gamma_{1}^{2}$ which is indicated by the yellow solid line. This is also predicted from the complex energy spectrum with OBC. The yellow dashed lines in Fig. 5(a) represent $m_{0}=2\pm\tilde{\gamma}$ lines as predicted from the complex energy under PBC. Interestingly, the real space invariant QWN fails to identify these phase boundaries. Unlike the Hermitian system, the phase boundaries between topological and trivial phases cannot be captured by the energy spectrum under OBC and PBC in the present NH case. This clearly suggests that the topological phase, predicted from the complex energy spectrum in OBC, is apprehended by the non-Bloch and bi-orthogonalized version of QWN. We note that the analytically derived phase boundaries are valid for $\gamma_{1,2}\ll 1$.
@@
We find qualitatively similar results in case of the second NN model as shown in Fig. 5(b). In addition to $N_{xy}=1$ phase, we obtain $N_{xy}=4$ phase where four zero-energy modes with ${\rm Re}[E]=0$ are present at each corner. While investigating the phase boundaries, it is expected to find $N_{xy}=1$ [$4$] for $3\gamma_{2}^{2}<m_{0}<4+5\gamma_{2}^{2}$ [$-2-\gamma^{2}_{2}<m_{0}<3\gamma_{2}^{2}$]. For smaller strength of non-Hermiticity i.e.,  $\gamma\to 0$, we find quantitative agreement between the analytical and numerical findings. The phase boundaries $m_{0}=4+5\gamma_{2}^{2}$, $3\gamma_{2}^{2}$, and $-2-\gamma^{2}_{2}$, designated by the yellow solid line, do not fully comply with the $N_{xy}$ profile for $\gamma>0.1$. This can be due to more intricacies than just the finite size effect. Interestingly, the following tendency is noticed: for $\tilde{\gamma}<m_{0}<4+5\gamma_{2}^{2}$, one obtains $N_{xy}=1$ while within the regime $-2-\tilde{\gamma}<m_{0}<-\tilde{\gamma}$, $N_{xy}$ acquires the value $4$. Therefore, the non-Bloch and bi-orthogonalized version of QWN can quantitatively and qualitatively identify the phase boundaries of SOTI phases hosting four and sixteen corner modes starting from the trivial phases for $m_{0}>4+5\gamma_{2}^{2}$ and $m_{0}<-2-\tilde{\gamma}$, respectively, across which $N_{xy}$ jumps between zero and finite values. The real space invariant QWN is thus a useful topological marker to identify the exceptional phases for our NH system with OBC.
On the other hand, between two SOTI phases with different number of corner modes (for $-\tilde{\gamma}<m_{0}<\tilde{\gamma}$), we find $N_{xy}\neq 0$ as depicted in Fig. 5(b). In the complex energy analysis under OBC, we find macroscopic degeneracies with ${\rm Re}[E]=0$ for $-2-\gamma^{2}_{2}<m_{0}<\tilde{\gamma}$. Likewise the earlier first NN case, $N_{xy}$ does not exhibit any jumps between finite and zero values around $m_{0}=4\pm\tilde{\gamma}$ and $m_{0}=-2\pm\tilde{\gamma}$ which are predicted by the complex energy spectrum under PBC. On the contrary, $m_{0}=\pm\tilde{\gamma}$ boundaries, predicted by complex energy spectrum under PBC, are visible as $N_{xy}$ changes between two finite values. The exceptional lines $m_{0}=-2\pm\tilde{\gamma}$, ±$\pm\tilde{\gamma}$, and $4\pm\tilde{\gamma}$ obtained employing PBC are depicted by yellow dashed lines in Fig. 5(b). This agreement is surprising and yet to be explored in the future. However, there is an apparent discrepancy between the solid yellow lines and the numerically obtained $N_{xy}$. This mismatch is owing to the fact that the mathematical form of non-Bloch transformation that we consider in the hopping terms while computing $N_{xy}$, employing PBC, is computed by employing a low-energy version of $H_{\gamma}(\boldsymbol{\mathrm{k}})$. To obtain the low-energy spectrum of $H_{\gamma}(\boldsymbol{\mathrm{k}})$, we expand the Hamiltonian around $\boldsymbol{\mathrm{k}}=(0,0)$. By doing that we can obtain the phase boundary associated with the right part of Fig. 5(b). However, when we incorporate the higher-order hopping elements, the low-energy model around $\boldsymbol{\mathrm{k}}=(0,0)$ does not necessarily encapsulate all the phase transition lines. In that scenarion, one should also consider a low-energy Hamiltonian around other momenta such as $\boldsymbol{\mathrm{k}}=(\pm 2\pi/3,\pm 2\pi/3)$, depending upon the value of $m_{0}$. Nevertheless, this scenario adds substantial complexity to the problem as one should consider a different non-Bloch form for different $m_{0}$. Thus, finding a universal transformation to obtain the exact phase boundary in the NH second NN hopping case still remains an interesting question and beyond the scope of the present paper. Nevertheless, complex energy spectra under PBC might be useful for understanding the phase boundaries between two different topological phases.
@@
Figure 5: The phase diagram in $m_{0}$-$\gamma$ plane for NH first NN and second NN hopping models are shown in panel (a), and (b), respectively. The color bar represents QWN $N_{xy}$. The phase boundary $m_{0}=2+\gamma_{1}^{2}$ [$m_{0}=2\pm\tilde{\gamma}$], obtained from OBC [PBC], between exceptional topological and trivial phases in panel (a) is identified by the yellow solid [dashed] line. The phase boundaries, associated with OBC [PBC], between topological phases hosting sixteen and four corner modes, and trivial phase are given by $m_{0}=4+5\gamma_{2}^{2}$ [$m_{0}=4\pm\tilde{\gamma}$], $m_{0}=3\gamma_{2}^{2}$ [$m_{0}=\pm\tilde{\gamma}$], and $m_{0}=-2-\gamma_{2}^{2}$ [$m_{0}=-2\pm\tilde{\gamma}$], respectively, represented by the yellow solid [dashed] lines in panel (b).

### v2

@@ III.1 NH Model with first NN hopping
We choose $m_{0}=1.0$, while the other model parameters remain the same as mentioned in Fig. 1.
@@ III.2 NH Model with second NN hopping
We now examine the energy spectrum for the second NN model in Figs. 3(d,e,f) employing OBC. Similar to the first NN case, the momentum takes the following non-Bloch form $k_{i}\to k^{\prime}_{i}-i\gamma_{2}$ with $\gamma_{2}=\gamma/(\lambda^{s}_{1}+2\lambda^{s}_{2})$ and $i=x,y$. This leads to the renormalized mass term $m^{\prime}_{0}=m_{0}-4-5\gamma_{2}^{2}<0~{}(>0)$ for the topological (trivial) phase with zero-energy (finite energy bulk) modes considering the Bloch momentum $\boldsymbol{\mathrm{k}}=(0,0)$. On the other hand, another topological (trivial) phase with zero-energy (finite energy) modes appears for $-m^{\prime}_{0}=m_{0}+2+\gamma_{2}^{2}>0~{}(<0)$ while exploiting energy around the Bloch momentum $\boldsymbol{\mathrm{k}}=(\pm 2\pi/3,\pm 2\pi/3)$. Our analysis indicates the existence of four [sixteen] corner modes for $3\gamma_{2}^{2}<m_{0}<4+5\gamma_{2}^{2}$ [$-2-\gamma^{2}_{2}<m_{0}<3\gamma_{2}^{2}$] yielding ${\rm Re}[E]=0$ under OBC. However, the numerical findings shown in Fig. 3(d) and the topological regime highlighted by the red lines do not fully match with the exceptional boundaries predicted analytically. This can be attributed to the finite size effect in the second NN case, which is substantially small for the first NN case. In particular, the rightmost boundary in Fig. 3(d) is more affected due to the finite size scaling as there exists a significant mismatch between the analytical prediction $m_{0}=4+5\gamma_{2}^{2}$ and real zero-energy modes obtained under OBC. To this end, we discuss the finite size scaling around the phase transition point at the rightmost part of Fig. 3(d) in Appendix B. Note that, the finite size effect is expected to become more substantial in case of OBC rather than for PBC. Due to this reason, we choose the phase boundary at $m_{0}=4+5\gamma_{2}^{2}$ for the finite size analysis. We find macroscopic degeneracies at ${\rm Re}[E]=0$ around $m_{0}=0$, unlike the previous case. The complex energy spectrum ${\rm Im}[E]$ does not manifest any noteworthy features, as shown in Figs. 3(b,e), irrespective of PBC and OBC cases. The absolute value of energy is expected to vanish i.e.,  $|E|=0$ under OBC for $-2-\gamma^{2}_{2}<m_{0}<4+5\gamma_{2}^{2}$. However, numerical results suffer from finite size effects as far as the exceptional boundaries are concerned [see the red lines in Fig. 3(f)]. The finite value of $\gamma$ (NH effect) thus extends the Hermitian topological phase beyond its boundaries, leading to exceptional topological phases.
@@ III.3 Quadrupole Winding number
We now examine the present situation with $\gamma\neq 0$ where the NH analog of QWN is discussed. Importantly, CS is also preserved for the NH Hamiltonian $CH_{\gamma}(\boldsymbol{\mathrm{k}})C^{-1}=-H_{\gamma}(\boldsymbol{\mathrm{k}})$ allowing for the anti-diagonal form of $\tilde{H}_{\gamma}$. At the same time, the definition of $U^{A,B}_{C}$ remains unaltered for the NH case. We adopt the bi-orthogonalized definitions of $U_{A}$ and $U^{\dagger}_{B}$, obtained from the singular value decomposition of $h$, to define $U^{\dagger}_{A}$ and $U_{B}$, respectively. One has to ensure $\sum_{n}|U^{R}_{\sigma,n}\rangle\langle U^{L}_{\sigma,n}|=\mathds{1}$ and $\langle U^{L}_{\sigma,n}|U^{R}_{\sigma,m}\rangle=\delta_{mn}$ with $\sigma=A,B$ and $L(R)$ denotes the left (right) singular vectors. This results in left [right] singular vectors corresponding to right singular vectors $(U^{\dagger}_{B})^{\dagger}\equiv U^{R}_{B}$ [left singular vectors $U^{\dagger}_{A}\equiv U^{L}_{A}$] as $U^{\dagger}_{B}\equiv U^{L}_{B}$ [$U_{A}\equiv U^{R}_{A}$]. Therefore, sublattice quadrupole operator $\mathcal{Q}_{\sigma}$ takes the form $\mathcal{Q}_{\sigma}=U_{\sigma}^{L}U^{\sigma}_{C}\mathcal{Q}U^{\sigma}_{C}U^{R}_{\sigma}$. In addition, the non-Bloch form of momentum has to be incorporated while computing $\mathcal{Q}_{\sigma}$.
@@
To be precise, the complex momentum $k_{i}\to k^{\prime}_{i}-i\gamma_{2}$ with $i=x,y$ leads to the exponentially enhanced and suppressed hopping elements by the multiplicative factors $\exp(\gamma_{1})$ and $\exp(-\gamma_{1})$ [$\exp(\gamma_{2})$ and $\exp(-\gamma_{2})$] for first [second] NN models. We use the real space form of the tight-binding model with the renormalized hopping amplitudes as follows $\lambda^{s,h,f}_{1,2}\to\lambda^{s,h,f}_{1,2}\exp(\gamma_{1,2})$ and $\lambda^{s,h,f}_{1,2}\to\lambda^{s,h,f}_{1,2}\exp(-\gamma_{1,2})$ for forward and backward hopping amplitudes, respectively. We consider the real space version of NH Hamiltonian $H_{\gamma}$ with the above-mentioned renormalized hoppings in order to compute QWN. Altogether, this enables us to define the NH analog of QWN $N_{xy}$ with dressed hopping and bi-orthogonalized definition. Note that the real part of $N_{xy}$ exhibits quantized value for the present case with $\gamma\neq 0$ as demonstrated below.
@@
We now discuss the phase diagram of the first and second NN NH model in the $\gamma$-$m_{0}$ plane in Figs. 5(a,b) respectively. Note that, there exists four corner modes with ${\rm Re}[E]=0$ for $m_{0}<2+\gamma_{1}^{2}$, yielding $N_{xy}=1$. This refers to the fact that there exists only one topological zero-mode per corner [see Fig. 5(a)]. On the other hand, when $m_{0}>2+\gamma_{1}^{2}$, the NH model does not host any topological phase and hence $N_{xy}=0$. Hence, the topological phase boundary is $m_{0}=2+\gamma_{1}^{2}$ which is indicated by the yellow solid line. This is also predicted from the complex energy spectrum with OBC. The yellow dashed lines in Fig. 5(a) represent $m_{0}=2\pm\tilde{\gamma}$ lines as predicted from the complex energy under PBC. Interestingly, the real space invariant QWN fails to identify these phase boundaries. Unlike the Hermitian system, the phase boundaries between topological and trivial phases cannot be captured by the energy spectrum under OBC and PBC in the present NH case. This clearly suggests that the topological phase, predicted from the complex energy spectrum in OBC, is apprehended by the non-Bloch and bi-orthogonalized version of QWN. We note that the analytically derived phase boundaries are valid for $\gamma_{1,2}\ll 1$. Interestingly, the non-Hermiticity induces additional regions $-2-\gamma_{1}^{2}<m_{0}<-2$ and $2<m_{0}<2+\gamma_{1}^{2}$ around $m_{0}=\pm 2$ beyond the Hermitian phase boundaries. In Fig. 5(a), we only illustrate the positive $m_{0}$ window where the NH SOTI phase is present for $m_{0}<2+\gamma_{1}^{2}$.
@@
We find qualitatively similar results in the case of the second NN model, as shown in Fig. 5(b). In addition to $N_{xy}=1$ phase, we obtain $N_{xy}=4$ phase where four zero-energy modes with ${\rm Re}[E]=0$ are present at each corner. While investigating the phase boundaries, it is expected to find $N_{xy}=1$ [$4$] for $3\gamma_{2}^{2}<m_{0}<4+5\gamma_{2}^{2}$ [$-2-\gamma^{2}_{2}<m_{0}<3\gamma_{2}^{2}$]. For smaller strength of non-Hermiticity i.e.,  $\gamma\to 0$, we find quantitative agreement between the analytical and numerical findings. The phase boundaries $m_{0}=4+5\gamma_{2}^{2}$, $3\gamma_{2}^{2}$, and $-2-\gamma^{2}_{2}$, designated by the yellow solid line, do not fully comply with the $N_{xy}$ profile for $\gamma>0.1$. This can be due to more intricacies than just the finite-size effect. Interestingly, the following tendency is noticed: for $\tilde{\gamma}<m_{0}<4+5\gamma_{2}^{2}$, one obtains $N_{xy}=1$ while within the regime $-2-\tilde{\gamma}<m_{0}<-\tilde{\gamma}$, $N_{xy}$ acquires the value $4$. Therefore, the non-Bloch and bi-orthogonalized version of QWN can quantitatively and qualitatively identify the phase boundaries of SOTI phases hosting four and sixteen corner modes starting from the trivial phases for $m_{0}>4+5\gamma_{2}^{2}$ and $m_{0}<-2-\tilde{\gamma}$, respectively, across which $N_{xy}$ jumps between zero and finite values. The real space invariant QWN is thus a useful topological marker to identify the exceptional phases for our NH system with OBC.
@@
Figure 5: The phase diagram in $m_{0}$-$\gamma$ plane for NH first NN and second NN hopping models are shown in panel (a), and (b), respectively. The color bar represents QWN $N_{xy}$. The phase boundary $m_{0}=2+\gamma_{1}^{2}$ [$m_{0}=2\pm\tilde{\gamma}$], obtained from OBC [PBC], between exceptional topological and trivial phases in panel (a) is identified by the yellow solid [dashed] line. The phase boundaries, associated with OBC [PBC], between topological phases hosting sixteen and four corner modes and trivial phase, are given by $m_{0}=4+5\gamma_{2}^{2}$ [$m_{0}=4\pm\tilde{\gamma}$], $m_{0}=3\gamma_{2}^{2}$ [$m_{0}=\pm\tilde{\gamma}$], and $m_{0}=-2-\gamma_{2}^{2}$ [$m_{0}=-2\pm\tilde{\gamma}$], respectively, represented by the yellow solid [dashed] lines in panel (b).
As mentioned earlier for the first NN case, the Hermitian phase boundaries are modified due to the non-Hermiticity. For second NN Hermitian counterpart with $\gamma=0$, one obtains SOTI phases $40<m_{0}<$ 4 and $-2<m_{0}<0$ hosting four and sixteen zero-energy corner modes, respectively. The NH factor $\gamma$ introduces four corner modes for positive values of $m_{0}$ beyond $m_{0}=4$ until $m_{0}<4+5\gamma_{2}^{2}$, as demonstrated in Fig. 5(b). The same applies to the negative values of $m_{0}$ where the sixteen NH corner modes continue to exist for $|m_{0}|<2+\tilde{\gamma}$ beyond $m_{0}=-2$. Importantly, in the second NN SOTI model, the number of corner modes changes for positive and negative values of $m_{0}$ which is not the case for the first NN SOTI model. Therefore, we would like to emphasize that the NH factor $\gamma\neq 0$ and the second NN hoppings $\lambda^{h}_{2},\lambda^{s}_{2},\lambda^{f}_{2}\neq 0$, together modify the phase diagram in a complex manner such that the number of corner modes and their corresponding parameter window vary significantly as compared to the first NN Hermitian model. It would be interesting to study in future why the changes in QWN do not always follow the OBC energy gap closing lines. For example, $m_{0}=-2-\gamma_{2}^{2}$ and $m_{0}=-2-\tilde{\gamma}$ ($m_{0}=3\gamma_{2}^{2}$ and $m_{0}=\pm\tilde{\gamma}$) phase boundaries around $m_{0}=-2$ ($m_{0}=0$) can be investigated further for better understanding of the interplay between NH term and second NN hoppings.
On the other hand, between two SOTI phases with different number of corner modes (for $-\tilde{\gamma}<m_{0}<\tilde{\gamma}$), we find $N_{xy}\neq 0$ as depicted in Fig. 5(b). In the complex energy analysis under OBC, we find macroscopic degeneracies with ${\rm Re}[E]=0$ for $-2-\gamma^{2}_{2}<m_{0}<\tilde{\gamma}$. Likewise, the earlier first NN case, $N_{xy}$ does not exhibit any jumps between finite and zero values around $m_{0}=4\pm\tilde{\gamma}$ and $m_{0}=-2\pm\tilde{\gamma}$ which are predicted by the complex energy spectrum under PBC. On the contrary, $m_{0}=\pm\tilde{\gamma}$ boundaries, predicted by complex energy spectrum under PBC, are visible as $N_{xy}$ changes between two finite values. The exceptional lines $m_{0}=-2\pm\tilde{\gamma}$, ±$\pm\tilde{\gamma}$, and $4\pm\tilde{\gamma}$ obtained employing PBC are depicted by yellow dashed lines in Fig. 5(b). This agreement is surprising and yet to be explored in the future. However, there is an apparent discrepancy between the solid yellow lines and the numerically obtained $N_{xy}$. This mismatch is owing to the fact that the mathematical form of non-Bloch transformation that we consider in the hopping terms while computing $N_{xy}$, employing PBC, is computed by employing a low-energy version of $H_{\gamma}(\boldsymbol{\mathrm{k}})$. To obtain the low-energy spectrum of $H_{\gamma}(\boldsymbol{\mathrm{k}})$, we expand the Hamiltonian around $\boldsymbol{\mathrm{k}}=(0,0)$. By doing that, we can obtain the phase boundary associated with the right part of Fig. 5(b). However, when we incorporate the second NN hopping elements, the low-energy model around $\boldsymbol{\mathrm{k}}=(0,0)$ does not necessarily encapsulate all the phase transition lines. In that scenario, one should also consider a low-energy Hamiltonian around other momenta such as $\boldsymbol{\mathrm{k}}=(\pm 2\pi/3,\pm 2\pi/3)$, depending upon the value of $m_{0}$. Nevertheless, this scenario adds substantial complexity to the problem as one should consider a different non-Bloch form for different $m_{0}$. Thus, finding a universal transformation to obtain the exact phase boundary in the NH second NN hopping case still remains an interesting question and is beyond the scope of the present paper. Nevertheless, complex energy spectra under PBC might be useful for understanding the phase boundaries between two different topological phases.
Symmetry Operation Remarks
Mirror-$x$ $M_{x}=\sigma_{0}s_{y}$: $M_{x}H_{\gamma}(k_{x},k_{y})M_{x}^{-1}=H_{\gamma}(-k_{x},k_{y})$ Broken, when $\lambda^{f}_{1,2},\gamma_{x}\neq 0$
Mirror-$y$ $M_{y}=\sigma_{z}s_{x}$: $M_{y}H_{\gamma}(k_{x},k_{y})M_{y}^{-1}=H_{\gamma}(k_{x},-k_{y})$ Broken, when $\lambda^{f}_{1,2},\gamma_{y}\neq 0$
Four-fold rotation $C_{4}=e^{-\frac{i\pi}{4}\sigma_{z}s_{z}}$: $C_{4}H_{\gamma}(k_{x},k_{y})C_{4}^{-1}=H_{\gamma}(-k_{y},k_{x})$ Broken, when $\lambda^{f}_{1,2},\gamma_{x,y}\neq 0$
Mirror-rotation I $M_{xy}=C_{4}M_{y}$: $M_{xy}H_{\gamma}(k_{x},k_{y})M_{xy}^{-1}=H_{\gamma}(k_{y},k_{x})$ Broken, when $\gamma_{x}\neq\gamma_{y}$
Mirror-rotation II $M_{x\bar{y}}=C_{4}M_{x}$: $M_{x\bar{y}}H_{\gamma}(k_{x},k_{y})M_{x\bar{y}}^{-1}=H_{\gamma}(-k_{y},-k_{x})$ Broken, when $\gamma_{x}\neq\gamma_{y}$
Table 1: Spatial symmetries and their operations are highlighted.